OpenStudy (anonymous):

(X+5)^2-12(X+5)+32=0 SOLVE THIS EQUATION BY USING THE QUADRATIC FORM

OpenStudy (tkhunny):

Well, do it! \((x+5) = What?\)

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

WHAT ABOUT THE REST OF IT

OpenStudy (tkhunny):

You have used the Quadratic Formula, right? What would you do if it said \(z^{2} - 12z + 32 = 0\)?

OpenStudy (anonymous):

I'D FACTOR RIGHT (Z-8)(Z-4)

OpenStudy (anonymous):

Z=8,4

OpenStudy (tkhunny):

I suppose you could do that. Are we supposed to use the Quadratic Formula? Anyway, your factoring is fine. Do this problem EXACTLY the same way. Just write "x+5" instead of "z".

OpenStudy (anonymous):

X+5-8 X+5-4

OpenStudy (anonymous):

X=-3,1?

OpenStudy (tkhunny):

You're almost done. Write "= 0" and solve. You missed some signs. Write THE ENTIRE equation before attempting to solve. (X+5) - 8 = 0 (X+5) - 4 = 0

OpenStudy (anonymous):

(X+5)=8 X=3? (X+5)=4 X=-1

OpenStudy (tkhunny):

Good work. It helps to write it ALL! :-)

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

(X-4)^2+4(X-4)-12=0 X^2+4X-12 (X+6)(X-2) X=-6,2 (X-4)-6=0 (X-4)+2=0 X-4=6...X=10? X-4=-2....X=2?

OpenStudy (tkhunny):

Pretty good, but it is confusing when you use the same variable name. You have X-4 and you changed to X. Better to change to Z or something other than X.